OpenStudy (anonymous):

Find the area of a rectangle if the perimeter is equal to 38 in and the length is 10 in.

OpenStudy (anonymous):

We are trying to find the width so we can work out the area Area=length*width=l*w It says l=10 and also that the perimeter is 38 Perimeter=2l+2w Solve for w and then find out the area :)

OpenStudy (anonymous):

how did you get 28 paremeter

OpenStudy (anonymous):

The question says the perimeter is 38

OpenStudy (anonymous):

sorry 21 perimeter

OpenStudy (anonymous):

Ohh that is 2*L+2*W=2L+2W

OpenStudy (anonymous):

when you said 21 + 2w

OpenStudy (anonymous):

oic

OpenStudy (anonymous):

seeme like i got 9 for width

OpenStudy (anonymous):

Correct!

OpenStudy (anonymous):

awesome ty

OpenStudy (anonymous):

Don't forget to work the area out :) No problem

OpenStudy (anonymous):

so length10 x width 9? then

OpenStudy (anonymous):

Yes A=L*W=10*9=90

OpenStudy (anonymous):

that's what i got too TY

OpenStudy (anonymous):

zed when you get a chance i have a math question for you

OpenStudy (anonymous):

@sharon penn are you given any other info e.g area or perimeter?

OpenStudy (anonymous):

@tinkerbell can you post the link?